OpenStudy (anonymous):

i need help -2 square root 45 - 3 square root 20 - 2 square root 6

OpenStudy (anonymous):

\[-2\sqrt{45}-3\sqrt{20}-2\sqrt{6}\]

OpenStudy (anonymous):

adding and subtracting radicals